In how many ways can the letters of the word ALLAHABAD be arranged so that the vowels can occupy only even positions.
Answer by sudhanshu_kmr(1152)

here 4 vowels and all are 'A'
here total 9 places.. 5 odds and 4 even places i.e OEOEOEOEO
no. of ways to arrange four 'A' at 4 places = 4P4/4! = 1


no. of ways to arrange 5 consonants where there are 2 'L' = 5!/2! = 60

total no. of ways = 1* 60 = 60

If any problem related to understand the concept, you are welcome to contact...